# Failure Modes of Maximum Entropy RLHF

### v1

@@ Abstract
In this paper, we show that Simple Preference Optimization (SimPO) can be derived as Maximum Entropy Reinforcement Learning with length-normalized temperature, providing a theoretical foundation for this reference-free method. Motivated by SimPO’s strong performance in offline preference optimization, we investigate whether Maximum Entropy RL can achieve similar results in online RLHF settings. Our experiments find that Maximum Entropy RL consistently exhibits overoptimization and unstable KL dynamics, even at very low learning rates. Unlike KL-constrained methods that maintain stable training, entropy regularization fails to prevent reward hacking and appears to correlate with overoptimization. Lastly, we discuss possible explanations for why SimPO succeeds in offline settings while Maximum Entropy RL struggles in online scenarios. Our findings suggest that reference-free approaches may face distinct challenges when applied to online or offline preference learning.
@@ 1 Introduction
Aligning AI systems with human values is widely recognized as a central challenge in modern AI (Bengio et al., 2025; Russell, 2022). The prevailing approach, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2023; Stiennon et al., 2022; Ziegler et al., 2020; Bai et al., 2022; Ouyang et al., 2022), typically follows a three-stage pipeline: (1) supervised fine-tuning (SFT), (2) training a reward model from preference data, and (3) optimizing the policy with reinforcement learning under KL divergence regularization to limit deviation from a reference model. While this framework has been successful, it is computationally demanding and operationally complex, requiring separate reward models, substantial human annotation, and careful hyperparameter tuning to balance reward maximization with stability.
@@
These limitations have motivated the exploration of direct alignment algorithms (DAAs) (Rafailov et al., 2024a) that aim to simplify the pipeline by avoiding explicit reward modeling. Direct Preference Optimization (DPO) (Rafailov et al., 2024c) is one such method, reformulating preference learning as a supervised objective with an implicit KL prior, and grounding its design in KL-constrained reinforcement learning. More recently, Simple Preference Optimization (SimPO) (Meng et al., 2024) has attracted attention for achieving strong empirical results while discarding the reference model entirely. Instead, SimPO employs length-normalized log likelihood and a target margin between preferred and dispreferred responses, yielding an objective that is simple to implement yet competitive in practice.
Despite these promising results, SimPO has lacked the kind of principled theoretical framework that underpins reference-based methods like DPO. This raises several questions: What might explain SimPO’s effectiveness as a reference-free approach? Can it be connected to established reinforcement learning principles? And if so, what might such a connection imply for the broader landscape of preference optimization methods?
@@
In this work, we take a step toward answering these questions by establishing a connection between SimPO and Maximum Entropy Reinforcement Learning (Ziebart et al., 2008). We show that SimPO can be interpreted as a closed-form solution to a Maximum Entropy RL objective with length-normalized temperature scaling. This perspective provides SimPO with a theoretical grounding analogous to DPO’s relationship with KL-constrained RL, while also suggesting that reference-free optimization may arise naturally from entropy regularization under certain conditions.
@@
At the same time, this analysis raises an empirical question: if SimPO can be viewed as an offline Maximum Entropy solution, could online Maximum Entropy RL also serve as a viable alternative to KL-constrained methods in RLHF? To explore this possibility, we conducted experiments comparing Maximum Entropy RL and KL-constrained RL on the TL;DR summarization benchmark using models from the Pythia suite.
@@
Our experiments reveal a notable asymmetry. While SimPO performs well in offline preference optimization, online Maximum Entropy RL often exhibited instability and signs of overoptimization, even at conservative learning rates. We also observed that increases in entropy tended to correlate with such instabilities, suggesting that entropy regularization may not always guard against reward hacking and, in some cases, could contribute to it. One possible explanation is that SimPO benefits from implicit stabilizing factors—such as dataset constraints and target margins—that approximate the regularization effects of a reference model, whereas these protections are absent in online Maximum Entropy RL.
@@
Our contributions are threefold. First, we provide a theoretical interpretation of SimPO as Maximum Entropy RL with adaptive temperature scaling, situating it within established RL frameworks. Second, we present empirical evidence that while Maximum Entropy RL is effective in offline settings (through SimPO), applying it directly in online RLHF can lead to instability and overoptimization, highlighting potential limitations of entropy regularization on its own. Third, we offer insight into why SimPO appears to succeed offline despite these challenges, pointing to the role of dataset constraints and target margins in stabilizing optimization. Together, these results provide a principled perspective on SimPO while suggesting that reference-free approaches may face important limitations in online training. We hope these findings help clarify the relationship between entropy-based methods and preference optimization, and open the door for further work on identifying the regularization mechanisms needed for robust online alignment.
@@ 2.1 Canonical RLHF
We reiterate the standart RLHF pipeline as outlined in (Ziegler et al., 2020) and subsequent works (Stiennon et al., 2022; Bai et al., 2022; Ouyang et al., 2022). It consists of three main stages: (1) Supervised Fine-Tuning (SFT), (2) Reward Modeling, and (3) RL Optimization.
@@
Prompts $x$ are sampled, and $\pi^{\text{SFT}}$ generates answer pairs $(y_{1},y_{2})$. Human annotators indicate preferences $y_{w}\succ y_{l}\mid x$, assumed to reflect a latent reward function $r^{*}(x,y)$. A common approach is to model preferences with the Bradley-Terry (BT) model (Bradley & Terry, 1952):
$p(y_{1}\succ y_{2}\mid x)=\frac{\exp\left(r^{*}(x,y_{1})\right)}{\exp\left(r^{*}(x,y_{1})\right)+\exp\left(r^{*}(x,y_{2})\right)}.$ (1)
@@
$\mathcal{L}R=-\mathbb{E}{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right],$ (2)
@@
$\max_{\pi_{\theta}}\mathbb{E}{x,y\sim\pi_{\theta}}[r_{\phi}(x,y)]-\beta\mathbb{D}_{\textrm{KL}}[\pi{{}_{\theta}}(y\mid x)||\pi_{\text{ref}}(y\mid x)].$ (3)
This prevents overoptimization and distributional shift. In practice, this objective is optimized with PPO (Schulman et al., 2017), using a reward defined as $r(x,y)=r_{\phi}(x,y)-\beta(\log\pi_{\theta}(y\mid x)-\log\pi_{\text{ref}}(y\mid x))$.
@@ 2.2 Direct Preference Optimization
Direct Preference Optimization (DPO) (Rafailov et al., 2024c) has become a popular method for preference-based tuning. Unlike traditional approaches that train a separate reward model, DPO defines the reward directly in terms of the optimized policy:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x),$ (4)
@@
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right],$ (5)
@@ 2.3 Simple Preference Optimization
Simple Preference Optimization (SimPO) (Meng et al., 2024) is a reference-free method for preference-based fine-tuning that aligns the reward used in training with the likelihood used at inference. Unlike DPO, SimPO eliminates the need for a reference policy by defining the reward as the length-normalized log-likelihood of the model output:
$r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)=\frac{\beta}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i})$ (6)
@@
$p(y_{w}\succ y_{l}\mid x)=\sigma\left(r(x,y_{w})-r(x,y_{l})-\gamma\right)$ (7)
@@
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}\mid x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}\mid x)-\gamma\right)\right]$ (8)
@@ 3 SimPO is the Maximum Entropy RL
SimPO is a widely used preference alignment method, appreciated for its strong empirical performance and simplicity due to its reference-free objective. However, it lacks a theoretical foundation, unlike reference-based approaches such as DPO, which is derived from a KL-constrained RL objective. Recent work (Liu et al., 2024) made the important observation that posterior probability rewards correspond to Maximum Entropy RL in their analysis of reference policies. Building on this insight, we establish the connection between this MaxEnt formulation and SimPO, showing that SimPO can be understood as Maximum Entropy RL with adaptive temperature through length normalization.
@@ 3.1 Maximum Entropy RL
Maximum Entropy Reinforcement Learning (MaxEnt RL) augments the standard RL objective with an entropy term, encouraging policies that align with the soft value function (Ziebart et al., 2008; Toussaint, 2009; Rawlik et al., 2013; Fox et al., 2015; O’Donoghue et al., 2016; Abdolmaleki et al., 2018; Haarnoja et al., 2018; Mazoure et al., 2020; Han & Sung, 2021; Zhang et al., 2025). It is deeply connected to probabilistic inference (Toussaint, 2009; Rawlik et al., 2013; Levine, 2018) and supported by both stochastic inference (Ziebart, 2010; Eysenbach & Levine, 2021) and game-theoretic foundations (Grünwald & Dawid, 2004; Ziebart et al., 2010; Han & Sung, 2021; Kim & Sung, 2023). MaxEnt is often favored for promoting exploration (Haarnoja et al., 2018; Hazan et al., 2019), smoothing optimization (Ahmed et al., 2019), and enabling robust decision-making (Eysenbach & Levine, 2021).
@@
$\pi^{\star}=\operatorname*{arg\,max}_{\pi}\;\mathbb{E}_{\tau\sim p^{\pi}(\tau)}\left[\sum_{t=1}^{T}r(\mathbf{s_{t}},\mathbf{a_{t}})+\alpha\,\mathcal{H}_{\pi}[\mathbf{a_{t}}\mid\mathbf{s_{t}}]\right],$ (9)
@@ 3.2 SimPO from Maximum Entropy RL
RLHF is commonly modeled as a contextual bandit problem, though some approaches treat it as a token-level MDP (Rafailov et al., 2024b; Xie et al., 2024). In this work, we adopt the contextual bandit view (Elwood et al., 2023), under which the maximum entropy formulation aligns with KL-constrained objectives. The resulting objective is given as follows.
@@
It is straightforward to show that optimal policy of the equation 10 (proof in Appendix  D) is as follows:
@@
$p^{*}(y_{1}\succ y_{2}|x)$ $=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)+\alpha\log Z(x)-\alpha\log\pi^{*}(y_{2}|x)-\alpha\log Z(x)\right)$ (14)
$=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)-\alpha\log\pi^{*}(y_{2}|x)\right)$ (15)
Crucially, the partition function $Z(x)$ cancels out, eliminating the need to compute it explicitly. To connect this to SimPO’s original formulation, we can decompose the temperature parameter $\alpha$ into two components: $\alpha=\frac{\beta}{|y|}$ where $\beta$ is a scaling factor and $|y|$ provides length normalization. Additionally, following ${\psi}PO$ (Azar et al., 2023), we can augment the objective with a target reward margin $\gamma>0$ to encourage separation between preferred and dispreferred responses. This leads to the SimPO objective for a parametric policy $\pi_{\theta}$:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right]$ (16)
This derivation reveals that SimPO is equivalent to Maximum Entropy RL under the contextual bandit formulation with adaptive temperature and target margin augmentation, making explicit the theoretical connection that underlies SimPO’s design. The reference-free nature of SimPO emerges naturally from the Maximum Entropy framework, as no explicit reference policy is required in the entropy-regularized objective. The length normalization can be interpreted as an adaptive temperature parameter that scales inversely with sequence length, while the target margin $\gamma$ encourages better separation between preferences.
@@
Following the same theoretical framework as DPO, SimPO inherits analogous guarantees regarding representational completeness, equivalence class preservation, and consistency under the Bradley-Terry preference model. The detailed proofs and formal statements of these properties are provided in Appendix D.

### v2

@@ Abstract
In this paper, we show that Simple Preference Optimization (SimPO) can be derived as Maximum Entropy Reinforcement Learning, providing a theoretical foundation for this reference-free method. Motivated by SimPO’s strong performance in offline preference optimization, we investigate whether Maximum Entropy RL can achieve similar results in online RLHF settings. Our experiments find that Maximum Entropy RL consistently exhibits overoptimization and unstable KL dynamics, even at very low learning rates. Unlike KL-constrained methods that maintain stable training, entropy regularization fails to prevent reward hacking and appears to correlate with overoptimization. Lastly, we discuss possible explanations for why SimPO succeeds in offline settings while Maximum Entropy RL struggles in online scenarios. Our findings suggest that reference-free approaches may face distinct challenges when applied to online or offline preference learning.
Failure Modes of Maximum Entropy RLHF
Ömer Veysel Çağatan KUIS AI Center, Koç University Sariyer/˙Istanbul,Turkey ocagatan19@ku.edu.tr          Barış Akgün Koç Universityi KUIS AI Center Sariyer/˙Istanbul,Turkey baakgun@ku.edu.tr
@@ 1 Introduction
Aligning AI systems with human values is widely recognized as a central challenge in modern AI (bengio2025singaporeconsensusglobalai; stuart). The dominant paradigm, Reinforcement Learning from Human Feedback (RLHF) (christiano2023deepreinforcementlearninghuman; stiennon2022learningsummarizehumanfeedback; ziegler2020finetuninglanguagemodelshuman; bai2022traininghelpfulharmlessassistant; ouyang2022traininglanguagemodelsfollow), typically follows a three-stage pipeline: supervised fine-tuning, reward model training from preference data, and policy optimization with reinforcement learning under KL divergence regularization to constrain deviation from a reference model. While effective, this pipeline is computationally expensive and operationally complex, requiring separate reward models, substantial human annotation, and careful hyperparameter tuning to maintain stability.
@@
These challenges have motivated direct alignment algorithms (DAAs) (rafailov2024scalinglawsrewardmodel), which bypass explicit reward modeling and online RL. Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) derives an analytical solution to a KL-regularized RL objective, expressing the reward implicitly as a function of the optimal policy and reducing preference learning to supervised optimization. More recently, Simple Preference Optimization (SimPO) (meng2024simposimplepreferenceoptimization) has demonstrated strong empirical performance while eliminating the reference model entirely, instead using length-normalized log-likelihoods and a fixed margin between preferred and dispreferred responses.
@@
Despite its empirical success, SimPO has lacked a principled theoretical foundation comparable to that of reference-based methods such as DPO. In this work, we establish a connection between SimPO and Maximum Entropy Reinforcement Learning (ziebart2008maximum). We show that SimPO can be interpreted as a closed-form solution to a Maximum Entropy RL objective, providing a theoretical grounding analogous to DPO’s relationship with KL-constrained RL and suggesting that reference-free optimization can naturally arise from entropy regularization under appropriate conditions.
@@
This perspective raises an important empirical question: if SimPO corresponds to an offline Maximum Entropy solution, can online Maximum Entropy RL serve as a viable alternative to KL-constrained methods in RLHF? To investigate this, we compare Maximum Entropy RL and KL-constrained RL on the TL;DR summarization benchmark (stiennon2022learningsummarizehumanfeedback) using models from the Pythia suite (biderman2023pythiasuiteanalyzinglarge).
@@
Our results reveal a clear asymmetry. While SimPO performs reliably in offline preference optimization, online Maximum Entropy RL frequently exhibits instability and overoptimization, even at conservative learning rates. We observe that increases in entropy often correlate with these instabilities, indicating that entropy regularization alone does not reliably prevent reward hacking and may, in some cases, exacerbate it. We hypothesize that SimPO benefits from implicit stabilizing mechanisms—such as dataset constraints and target margins—that partially substitute for the regularization provided by a reference model, whereas these protections are absent in online Maximum Entropy RL.
@@
Our contributions are threefold. First, we provide a theoretical interpretation of SimPO as a Maximum Entropy Reinforcement Learning solution, situating it within established RL frameworks. Second, we empirically demonstrate that directly applying Maximum Entropy RL in online RLHF settings can lead to instability and overoptimization, highlighting limitations of entropy regularization in isolation. Third, we offer insight into why SimPO succeeds offline despite these challenges, emphasizing the role of implicit regularization through data constraints and margin-based objectives. Together, these findings clarify the relationship between entropy-based methods and preference optimization, and point to the need for additional regularization mechanisms for robust reference-free alignment in online settings.
@@ 2.1 Canonical RLHF
We reiterate the standart RLHF pipeline as outlined in (ziegler2020finetuninglanguagemodelshuman) and subsequent works (stiennon2022learningsummarizehumanfeedback; bai2022traininghelpfulharmlessassistant; ouyang2022traininglanguagemodelsfollow). It consists of three main stages: (1) Supervised Fine-Tuning (SFT), (2) Reward Modeling, and (3) RL Optimization.
@@
Prompts $x$ are sampled, and $\pi^{\text{SFT}}$ generates answer pairs $(y_{1},y_{2})$. Human annotators indicate preferences $y_{w}\succ y_{l}\mid x$, assumed to reflect a latent reward function $r(x,y)$. A common approach is to model preferences with the Bradley-Terry (BT) model (19ff28b9-64f9-3656-ba40-08326a05748e):
$p(y_{1}\succ y_{2}\mid x)=\frac{e^{r^{*}(x,y_{1})}}{e^{r^{*}(x,y_{1})}+e^{r^{*}(x,y_{2})}}$ (1)
@@
$\mathcal{L}R$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\quad\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right]$
@@
$\mathcal{L}_{R}=-\mathbb{E}{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right],$
This prevents overoptimization and distributional shift. In practice, this objective is optimized with PPO (schulman2017proximalpolicyoptimizationalgorithms), using a reward defined as $r(x,y)=r_{\phi}(x,y)-\beta(\log(\pi_{\theta}(y\mid x))-\log(\pi_{\text{ref}}(y\mid x)))$.
@@ 2.2 Direct Preference Optimization
Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) has become a popular method for preference-based tuning. Unlike traditional approaches that train a separate reward model, DPO defines the reward directly in terms of the optimized policy:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x),$
@@
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right]$
@@ 2.3 Simple Preference Optimization
Simple Preference Optimization (SimPO) (meng2024simposimplepreferenceoptimization) is a reference-free method for preference-based fine-tuning that aligns the reward used in training with the likelihood used at inference. Unlike DPO, SimPO eliminates the need for a reference policy by defining the reward as the length-normalized log-likelihood of the model output:
$r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)=$
$\frac{\beta}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i})$
@@
$p(y_{w}\succ y_{l}\mid x)=\sigma\left(r(x,y_{w})-r(x,y_{l})-\gamma\right)$
@@
[…]
$\left.\left.\quad-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}\mid x)-\gamma\right)\right]$
@@ 3 SimPO is the Maximum Entropy RL
SimPO is a widely used preference alignment method, appreciated for its strong empirical performance and simplicity due to its reference-free objective. However, it lacks a theoretical foundation, unlike reference-based approaches such as DPO, which is derived from a KL-constrained RL objective. Recent work (liu2024understandingreferencepoliciesdirect) made the important observation that posterior probability rewards correspond to Maximum Entropy RL in their analysis of reference policies. Building on this insight, we establish the connection between this Maximum Entropy RL formulation and SimPO, showing that SimPO can be understood as Maximum Entropy RL with the addition of length normalization and target margins from preference learning.
@@ 3.1 Maximum Entropy RL
Maximum Entropy Reinforcement Learning (MaxEnt RL) augments the standard RL objective with an entropy term, encouraging policies that align with the soft value function (ziebart2008maximum; toussaint2009robot; rawlik2013stochastic; fox2015taming; o2016combining; abdolmaleki2018maximum; sac; mazoure2020leveraging; han2021max; zhang2025maximumentropymisleadspolicy). It is deeply connected to probabilistic inference (toussaint2009robot; rawlik2013stochastic; levine2018reinforcement) and supported by both stochastic inference (ziebart2010modeling; eysenbach2021maximum) and game-theoretic foundations (grunwald2004game; ziebart2010maximum; han2021max; kim2023adaptive). MaxEnt is often favored for promoting exploration (sac; hazan2019provably), smoothing optimization (ahmed2019understanding), and enabling robust decision-making (eysenbach2021maximum).
@@
$\pi^{\star}=\operatorname*{arg\,max}_{\pi}\;\mathbb{E}_{\tau\sim p^{\pi}(\tau)}$
$\left[\sum_{t=1}^{T}r(\mathbf{s_{t}},\mathbf{a_{t}})+\alpha\,\mathcal{H}_{\pi}[\mathbf{a_{t}}\mid\mathbf{s_{t}}]\right]$
@@ 3.2 SimPO from Maximum Entropy RL
RLHF is commonly modeled as a contextual bandit problem, though some approaches treat it as a token-level MDP (rafailov2024rqlanguagemodel; xie2024exploratorypreferenceoptimizationharnessing). In this work, we adopt the contextual bandit view (Elwood_2023), under which the maximum entropy formulation aligns with KL-constrained objectives. The resulting objective is given as follows.
@@
It is straightforward to show that optimal policy of the equation 2 (proof in Appendix  C) is as follows:
@@
$p^{*}(y_{1}$ $\succ y_{2}|x)$
$=\sigma\big(\alpha\log\pi^{*}(y_{1}|x)+\alpha\log Z(x)$
$\qquad-\alpha\log\pi^{*}(y_{2}|x)-\alpha\log Z(x)\big)$
$=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)-\alpha\log\pi^{*}(y_{2}|x)\right)$
Crucially, the partition function $Z(x)$ cancels out, eliminating the need to compute it explicitly. This establishes a direct connection between the Bradley–Terry preference model and the optimal policy induced by Maximum Entropy RL. To recover the SimPO training objective used in practice, we introduce two additional components that are motivated by empirical and structural considerations in preference learning rather than by the Maximum Entropy RL derivation itself.
First, we apply length normalization by scaling the temperature with the response length. This is necessary in SimPO because, unlike KL-constrained methods such as DPO where the reference policy induces implicit regularization through cancellation of log-probabilities, SimPO optimizes raw log-likelihoods and is therefore prone to severe length exploitation without explicit normalization. Second, inspired by ${\psi}PO$ (azar2023generaltheoreticalparadigmunderstand), we incorporate a target reward margin $\gamma>0$ as a fixed substitute for the adaptive margin implicitly provided by the reference policy in DPO. These additions shape the optimization landscape and improve empirical stability, but do not alter the underlying connection between SimPO and Maximum Entropy RL established above. This leads to the SimPO objective for a parametric policy $\pi_{\theta}$:
$\mathcal{L}_{\text{SimPO}}$ $(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg[\log\sigma\bigg($
$\quad\frac{\alpha}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)$
$\quad-\frac{\alpha}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\bigg)\bigg]$
This derivation reveals that SimPO is equivalent to Maximum Entropy RL under the contextual bandit formulation with addition of length normalization and target margin augmentation, making explicit the theoretical connection that underlies SimPO’s design.
@@
Following the same theoretical framework as DPO, SimPO inherits analogous guarantees regarding representational completeness, equivalence class preservation, and consistency under the Bradley-Terry preference model. The detailed proofs and formal statements of these properties are provided in Appendix C.